# The non-thermal emission following GW170817 is consistent with a conical radially-stratified outflow with initial Lorentz factor $\lesssim 10$

### v1

@@ 6 Conclusions
We have shown that the non-thermal radio to X-ray emission following the neutron star merger GW170817 is consistent with synchrotron emission from a collisionless shock driven into the interstellar medium (ISM) by a conical radially stratified outflow observed $\approx 0.25$ rad off-axis, with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-5}$ over the range $5\lesssim\gamma\lesssim 10$, opening (half-)angle $\approx 0.2$ rad, and total energy of $\approx 10^{51}$erg. A comparison of the measured and model-predicted fluxes and light centroid motion is given in figures 3 and 4, for a specific choice of values of model parameters. We did not attempt to obtain a "best-fit" set of parameter values, as the real outflow structure is likely more complicated than the simplified structure that we are using for our model calculations, which also includes simplifying assumptions regarding the microphysics of particle acceleration and emission of radiation (e.g. fixed electron and magnetic field energy fractions, independent of shock parameters). The results of the 2D numeric calculations of the light curve and centroid motion agree with those of our semi-analytic approximation (valid up to the flux peak time, see figure 2), supporting the validity of both.
@@
The radially stratified outflow model predicts a dependence on the observing angle, which is different than that predicted by highly relativistic jet models. Particularly, similar merger events observed at smaller viewing angles are predicted in our model to show a similarly extended duration of flux increase with time. Our analysis demonstrates that the data do not require a highly relativistic $\gamma\approx 100$ component, but the presence of such a component with opening angle $\ll 0.2$ rad and energy $\ll 10^{51}$ erg cannot be excluded, and it may dominate the observed flux when observed on-axis. The rapid deceleration of such a component, compared to the deceleration of the slower $\gamma<10$ outflow, will lead to a light curve that can be discriminated from that of the slower outflow.
Finally, we note that since flow structures that differ widely ($\gamma\sim 100$ relativistic jet and $\gamma<10$ radially stratified outflow) may predict similar non-thermal flux when observed off-axis, one cannot accurately infer parameters, e.g. the viewing and opening angles, based on fitting models to the data. This also implies that an accurate determination of $H_{0}$ (Hotokezaka et al., 2019; Govreen-Segal & Nakar, 2023; Palmese et al., 2024) is impossible through such parameter inference.
Acknowledgements

### v2

@@ 6 Conclusions
We have shown that the non-thermal radio to X-ray emission following the neutron star merger GW170817 is consistent with synchrotron emission from a collisionless shock driven into the interstellar medium (ISM) by a conical radially stratified outflow observed $\approx 0.25$ rad off-axis, with a power-low mass dependence on momentum, $M(>\gamma\beta)\propto(\gamma\beta)^{-4}$ over the range $5.5\lesssim\gamma\lesssim 10$, opening (half-)angle $\approx 0.15$ rad, and total energy of $\approx 5\times 10^{50}$erg. A comparison of the measured and model-predicted fluxes and light centroid motion is given in figures 3 and 5, for a specific choice of values of model parameters. We did not attempt to obtain a "best-fit" set of parameter values, as the real outflow structure is likely more complicated than the simplified structure that we are using for our model calculations, which also includes simplifying assumptions regarding the microphysics of particle acceleration and emission of radiation (e.g., fixed electron and magnetic field energy fractions, independent of shock parameters). The results of the 2D numeric calculations of the light curve and centroid motion agree with those of our semi-analytic approximation (valid up to the flux peak time, see figure 2), supporting the validity of both.
@@
The radially stratified outflow model predicts a dependence on the observing angle, which is different than that predicted by highly relativistic jet models. Particularly, similar merger events observed at smaller viewing angles are predicted in our model to show a similarly extended duration of flux increase with time. Our analysis demonstrates that the data do not require a highly relativistic $\gamma\approx 100$ component, but the presence of such a component with opening angle $\ll 0.15$ rad and energy $\ll 5\times 10^{50}$ erg cannot be excluded, and it may dominate the observed flux when observed on-axis (although slower outflows for gamma-ray bursts afterglows were suggested in Dereli-Bégué et al., 2022). The rapid deceleration of such a component, compared to the deceleration of the slower $\gamma<10$ outflow, will lead to a light curve that can be discriminated from that of the slower outflow. We note that since flow structures that differ widely ($\gamma\sim 100$ relativistic jet and $\gamma<10$ radially stratified outflow) may predict similar non-thermal flux when observed off-axis, (with relativistic jets typically predicting a steeper decline), one cannot accurately infer parameters, e.g., the viewing and opening angles, based on fitting models to the data. This also implies that an accurate determination of $H_{0}$ (Hotokezaka et al., 2019; Govreen-Segal & Nakar, 2023; Palmese et al., 2024) is impossible through such parameter inference.
Acknowledgments
@@
Data Availability
The data underlying this article will be shared following a reasonable request to the corresponding author.